# Rethinking Probabilistic Circuit Parameter Learning

### v1

@@ Abstract
Probabilistic Circuits (PCs) offer a computationally scalable framework for generative modeling, supporting exact and efficient inference of a wide range of probabilistic queries. While recent advances have significantly improved the expressiveness and scalability of PCs, effectively training their parameters remains a challenge. In particular, a widely used optimization method, full-batch Expectation-Maximization (EM), requires processing the entire dataset before performing a single update, making it ineffective for large datasets. While empirical extensions to the mini-batch setting have been proposed, it remains unclear what objective these algorithms are optimizing, making it difficult to assess their theoretical soundness. This paper bridges the gap by establishing a novel connection between the general EM objective and the standard full-batch EM algorithm. Building on this, we derive a theoretically grounded generalization to the mini-batch setting and demonstrate its effectiveness through preliminary empirical results.
@@ 1 Introduction
Probabilistic Circuits (PCs) are a class of generative models that represent probability distributions by recursively composing simpler distributions through sum (mixture) and product (factorization) operations [Choi et al., 2020]. The key idea behind PCs is to examine how tractable probabilistic models, such as Hidden Markov Models [Rabiner and Juang, 1986], perform inference (e.g., computing marginal probabilities). PCs distill the structure of these models’ computation graphs into a compact and general framework, which leads to a unified, computation-oriented perspective on tractable probabilistic modeling.
While significant progress has been made in improving the expressiveness of PCs through architectural innovations [Loconte et al., 2025, Liu and Van den Broeck, 2021] and system-level advancements [Liu et al., 2024, Peharz et al., 2020], there is still no clear consensus on how to effectively learn their parameters. Full-batch Expectation-Maximization (EM) and its empirical variants remain widely used approaches [Zhang et al., 2025, Liu et al., 2023]. However, full-batch EM requires aggregating information across the entire dataset before each parameter update, making it hard to scale to large datasets or streaming settings.
Based on Kunstner et al. [2021], which studies the full-batch EM algorithm for all latent variable models, we discover that the full-batch EM update of PCs corresponds to optimizing a 1st order Taylor approximation, regularized by a Kullback–Leibler (KL) divergence that penalizes deviation from the current distribution. This yields a novel view of the full-batch EM update for PCs, which has appeared in various forms across different contexts [Peharz, 2015, Choi et al., 2021, Poon and Domingos, 2011].
This perspective naturally suggests a theoretically grounded mini-batch extension: by increasing the weight on the KL term, we can compensate for the reduced information available in a mini-batch compared to the full dataset. The resulting update rule admits a closed-form expression, making it efficient and easy to implement. Preliminary empirical evaluations further demonstrate the empirical superiority compared to existing EM- and gradient-based algorithms.
@@ 2.1 Distributions as Circuits
Probabilistic Circuits (PCs) represent probability distributions with deep and structured computation graphs that consist of sum and product operations [Choi et al., 2020]. They are an extension of existing tractable probabilistic models, which are designed to support efficient and exact probabilistic inference over complex queries, such as Sum Product Networks [Poon and Domingos, 2011], cutset networks [Rahman et al., 2014], and Hidden Markov Models [Rabiner and Juang, 1986]. PCs inherit the tractability of these models while introducing a scalable computational framework to encode richer and more complex data distributions. The syntax and semantics of PCs are as follows:
@@
A PC ${p}$ over variables $\mathbf{X}$ is a directed acyclic computation graph with one single root node $n_{\mathrm{r}}$. Every input node (those without incoming edges) in ${p}$ defines an univariate distribution over variable $X\in\mathbf{X}$. Every inner node (those with incoming edges) is either a product or a sum node, where each product node encodes a factorized distribution over its child distributions and each sum node represents a weighted mixture of its child distributions. Formally, the distribution ${p}_{n}$ encoded by a node $n$ can be represented recursively as
$\!\!\!\!{p}_{n}(\boldsymbol{x})\!:=\!\begin{cases}f_{n}(\boldsymbol{x})&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ an\leavevmode\nobreak\ input\leavevmode\nobreak\ node},\\ \prod_{c\in\mathsf{ch}(n)}{p}_{c}(\boldsymbol{x})&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ product\leavevmode\nobreak\ node},\!\\ \sum_{c\in\mathsf{ch}(n)}\theta_{n,c}\cdot{p}_{c}(\boldsymbol{x})\!\!\!&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ sum\leavevmode\nobreak\ node},\end{cases}$ (1)
where $f_{n}$ is an univariate primitive distribution defined over $X\in\mathbf{X}$ (e.g., Gaussian, Categorical), $\mathsf{ch}(n)$ denotes the set of child nodes of $n$, and $\theta_{n,c}\geq 0$ is the parameter corresponds to the edge $(n,c)$ in the PC. Define the log-parameter of $(n,c)$ as $\phi_{n,c}:=\log\theta_{n,c}$, which will be used interchangeably with $\theta_{n,c}$. We further denote $\boldsymbol{\phi}:=\{\phi_{n,c}\}_{(n,c)}$ as the set of all sum node parameters in the PC. Without loss of generality, we assume that every path from the root node to an input node alternates between sum and product nodes.
In order to ensure exact and efficient computation of various probabilistic queries, including marginalization and computing moments, we need to impose structural constraints on the circuit. Specifically, smoothness and decomposability [Peharz et al., 2015] are a set of sufficient conditions that ensure tractable computation of marginal and conditional probabilities. Intuitively, this tractability arises because smooth and decomposable circuits represent multilinear functions, which are known to support efficient marginalization [Broadrick et al., 2024]. We provide further details in Appendix A.
PCs can be viewed as latent variable models with discrete latent spaces [Peharz et al., 2016]. Each sum node can be interpreted as introducing a discrete latent variable $Z$ that selects among its child distributions. Specifically, assigning $Z=i$ corresponds to choosing the $i$-th child of the sum node. By aggregating all such latent variables, the PC can be seen as defining a hierarchical latent variable model.
@@ 2.2 Expectation-Maximization
Expectation-Maximization (EM) is a well-known algorithm to maximize the log-likelihood given data $\boldsymbol{x}$ of a distribution defined by a latent variable model. Specifically, the distribution ${p}_{\boldsymbol{\phi}}(\mathbf{X})$ ($\boldsymbol{\phi}$ are the parameters) is defined as $\sum_{\boldsymbol{z}}{p}_{\boldsymbol{\phi}}(\mathbf{X},\boldsymbol{z})$ over latents $\mathbf{Z}$. Our goal is to maximize
@@
EM is an effective way to maximize the above objective when ${p}_{\boldsymbol{\phi}}(\mathbf{X},\mathbf{Z})$ permits much simpler (or even closed-form) maximum likelihood estimation. It optimizes Equation 2 by maximizing the following surrogate objective:
@@
Given the current set of parameters $\boldsymbol{\phi}$, EM updates the parameters by solving for $\boldsymbol{\phi}^{\prime}$ that maximizes $Q_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$, which is guaranteed to be a lower bound of $\mathtt{LL}(\boldsymbol{\phi}^{\prime})$ since
@@ 3 EM for Probabilistic Circuits
While variants of the EM algorithm have been proposed for training PCs in various contexts [Poon and Domingos, 2011, Peharz, 2015], their connection to the general EM objective $Q_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$ (cf. Eq. (3)) remains implicit. The lack of a unified formulation makes it difficult to fully understand the existing optimization procedures or to extend them to new settings, such as training with mini-batches of data, which is critical for scaling the optimizer to large datasets.
@@ 3.1 Revisiting Full-Batch EM
Given a PC ${p}_{\boldsymbol{\phi}}$ with log-parameters $\boldsymbol{\phi}$ (cf. Def. 1) and a dataset $\mathcal{D}$, the objective $Q_{\boldsymbol{\phi}}^{\mathcal{D}}(\boldsymbol{\phi}^{\prime})$ equals the following up to a constant term irrelevant to $\boldsymbol{\phi}^{\prime}$:
$\frac{1}{|\mathcal{D}|}\sum_{\boldsymbol{x}\in\mathcal{D}}\underbrace{\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})+\left\langle\frac{\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})}{\partial\boldsymbol{\phi}},\boldsymbol{\phi}^{\prime}-\boldsymbol{\phi}\right\rangle}_{\mathtt{LinLL}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(\boldsymbol{\phi}^{\prime})}-\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime}),$
@@
The proof follows Kunstner et al. [2021] and is provided in Section B.1. Proposition 1 reveals that the EM update can be interpreted as maximizing a regularized first-order approximation of the log-likelihood. Specifically, the term $\mathtt{LinLL}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(\boldsymbol{\phi}^{\prime})$ corresponds to the linearization of $\log{p}_{\boldsymbol{\phi}^{\prime}}(\boldsymbol{x})$ around the current parameters $\boldsymbol{\phi}$, capturing the local sensitivity of the log-likelihood to parameter changes. The KL term, $\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$, acts as a regularizer that penalizes large deviations in the joint distribution over $\mathbf{X}$ and $\mathbf{Z}$.
According to Proposition 1, solving for the updated parameters $\boldsymbol{\phi}^{\prime}$ requires computing two key quantities: (i) the gradient of the log-likelihood $\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})/\partial\boldsymbol{\phi}$, and the KL divergence $\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$. To express these terms in closed form, we introduce the concept of top-down probabilities, which is first defined by Dang et al. [2022].
@@
Given a PC ${p}$ parameterized by $\boldsymbol{\phi}$, we define the top-down probability (TD-prob) $\mathtt{TD}(n)$ of a node $n$ recursively from the root node to input nodes:
$\mathtt{TD}(n):=\begin{cases}1&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ the\leavevmode\nobreak\ root\leavevmode\nobreak\ node},\\ \sum_{m\in\mathsf{pa}(n)}\mathtt{TD}(m)&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ sum\leavevmode\nobreak\ node},\\ \sum_{m\in\mathsf{pa}(n)}\theta_{m,n}\cdot\mathtt{TD}(m)\!\!&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ product\leavevmode\nobreak\ node},\end{cases}$
@@
Assume the distributions defined by all nodes in a PC are normalized. We have the following ($\forall\boldsymbol{x}$):
(i) $\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})/\partial\boldsymbol{\phi}=\partial\log\hat{{p}}_{\boldsymbol{\phi}}(\boldsymbol{x})/\partial\boldsymbol{\phi}-\mathtt{TD}(\boldsymbol{\phi}),$
@@
The assumption that the PC is normalized is mild and practical. In Appendix C, we introduce a simple and efficient algorithm that adjusts the PC parameters to ensure normalization without affecting the structure of the circuit. We can now substitute the closed-form expressions for the gradient and the KL divergence into the general EM objective $Q_{\boldsymbol{\phi}}^{\mathcal{D}}(\boldsymbol{\phi}^{\prime})$, which converts the problem into [^3] We drop all terms that are independent of $\boldsymbol{\phi}^{\prime}$.
[…]
@@
where we define $\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c):=\frac{1}{|\mathcal{D}|}\sum_{\boldsymbol{x}\in\mathcal{D}}\frac{\partial\log\tilde{{p}}_{\boldsymbol{\phi}}(\boldsymbol{x})}{\partial\phi_{n,c}}$ [^4] This corresponds to the PC flows defined by Choi et al. [2021]., and $Z=\sum_{c^{\prime}\in\mathsf{ch}(n)}\mathtt{F}_{\boldsymbol{\phi}}^{\mathcal{D}}(n,c^{\prime})$ ensures that $n$ is normalized.
While this full-batch EM algorithm in Equation 4 has been derived in prior work [Choi et al., 2021, Peharz, 2015], we recover it here through Proposition 1. This paves the way for a principled mini-batch EM algorithm by generalizing the objective $Q_{\boldsymbol{\phi}}^{\mathcal{D}}(\boldsymbol{\phi}^{\prime})$, as shown in the next section.

### v2

@@ Abstract
Probabilistic Circuits (PCs) offer a computationally scalable framework for generative modeling, supporting exact and efficient inference of a wide range of probabilistic queries. While recent advances have significantly improved the expressiveness and scalability of PCs, effectively training their parameters remains a challenge. In particular, a widely used optimization method, full-batch Expectation-Maximization (EM), requires processing the entire dataset before performing a single update, making it ineffective for large datasets. Although empirical extensions to the mini-batch setting, as well as gradient-based mini-batch algorithms, converge faster than full-batch EM, they generally underperform in terms of final likelihood. We investigate this gap by establishing a novel theoretical connection between these practical algorithms and the general EM objective. Our analysis reveals a fundamental issue that existing mini-batch EM and gradient-based methods fail to properly regularize distribution changes, causing each update to effectively “overfit” the current mini-batch. Motivated by this insight, we introduce anemone, a new mini-batch EM algorithm for PCs. Anemone applies an implicit adaptive learning rate to each parameter, scaled by how much it contributes to the likelihood of the current batch. Across extensive experiments on language, image, and DNA datasets, anemone consistently outperforms existing optimizers in both convergence speed and final performance.
@@ 1 Introduction
Probabilistic Circuits (PCs) are a class of generative models that represent probability distributions by recursively composing simpler distributions through sum (mixture) and product (factorization) operations (Choi et al.,, 2020). The key idea behind PCs is to examine how tractable probabilistic models, such as Hidden Markov Models (Rabiner and Juang,, 1986), perform inference (e.g., computing marginal probabilities). PCs distill the structure of these models’ computation graphs into a compact and general framework, which leads to a unified, computation-oriented perspective on tractable probabilistic modeling.
While significant progress has been made in improving the expressiveness of PCs through architectural innovations (Loconte et al.,, 2025; Liu and Van den Broeck,, 2021) and system-level advancements (Liu et al.,, 2024; Peharz et al.,, 2020), there is still no clear consensus on how to effectively learn their parameters. Full-batch Expectation-Maximization (EM) and its empirical variants remain widely used approaches (Zhang et al.,, 2025; Liu et al., 2023a, ). However, full-batch EM requires aggregating information across the entire dataset before each parameter update, making it hard to scale to large datasets or streaming settings. Although mini-batch extensions and gradient-based optimization methods can converge faster, they typically achieve lower final log-likelihood than full-batch EM.
Based on Kunstner et al., (2021), which studies the full-batch EM algorithm for exponential-family latent variable models, we discover that the full-batch EM update of PCs corresponds to optimizing a 1st order Taylor approximation, regularized by a Kullback–Leibler (KL) divergence that penalizes deviation from the current distribution. This yields a novel view of the full-batch EM update for PCs, which has appeared in various forms across different contexts (Peharz,, 2015; Choi et al.,, 2021; Poon and Domingos,, 2011).
This perspective naturally suggests a theoretically grounded mini-batch extension: by increasing the weight on the KL term, we can compensate for the reduced information available in a mini-batch compared to the full dataset. The resultant algorithm, anemone (“an EM one”), adaptively applies large learning rates only to the PC parameters that strongly influence the likelihood of the current batch. The other parameters are kept (almost) unchanged to prevent excessive distribution drift, thereby ensuring that updates remain consistent with the data distribution.
The theoretical insights also explain the inferior performance of existing EM- and gradient-based mini-batch algorithms, as they regularize the distribution shift before and after an update using the KL divergence of the local distributions defined at sum nodes [^1] See Section 2 for the definition of sum nodes. and the L2 distance in the parameter space, respectively. As a result, these methods fail to effectively minimize the distribution shift while learning to improve the likelihood given the current batch, causing each update to “overfit” to the current samples and impede overall convergence.
Anemone admits a closed-form expression, making it efficient and easy to implement. We conduct extensive empirical evaluations on three types of datasets (language, image, and DNA) and four widely used classes of PC architectures. The results demonstrate that anemone consistently and significantly outperforms existing optimizers in both convergence speed and final likelihood.
@@ 2.1 Distributions as Circuits
Probabilistic Circuits (PCs) represent probability distributions with deep and structured computation graphs that consist of sum and product operations (Choi et al.,, 2020). They serve as a general framework encompassing tractable probabilistic models, which are designed to support efficient and exact probabilistic inference over complex queries, such as Sum Product Networks (Poon and Domingos,, 2011), cutset networks (Rahman et al.,, 2014), Hidden Markov Models (Rabiner and Juang,, 1986), and Probabilistic Generating Circuits (Zhang et al.,, 2021). The syntax and semantics of PCs are as follows:
@@
A PC p{p} over variables $\mathbf{X}$ is a directed acyclic computation graph with one single root node $n_{\mathrm{r}}$. Every input node (those without incoming edges) in p{p} defines an univariate distribution over variable $X\in\mathbf{X}$. Every inner node (those with incoming edges) is either a product or a sum node, where each product node encodes a factorized distribution over its child distributions and each sum node represents a weighted mixture of its child distributions. Formally, the distribution ${p}_{n}$ encoded by a node $n$ can be represented recursively as
[…]
where $f_{n}$ is an univariate primitive distribution defined over $X\in\mathbf{X}$ (e.g., Gaussian, Categorical), $\mathsf{ch}(n)$ denotes the set of child nodes of $n$, and $\theta_{n,c}\geq 0$ is the parameter corresponds to the edge $(n,c)$ in the PC. Define the log-parameter of $(n,c)$ as $\phi_{n,c}:=\log\theta_{n,c}$, which will be used interchangeably with $\theta_{n,c}$. We further denote $\boldsymbol{\phi}:=\{\phi_{n,c}\}_{(n,c)}$ as the set of all sum node parameters in the PC. Without loss of generality, we assume that every path from the root node to an input node alternates between sum and product nodes. [^2] This can be efficiently enforced since we can directly “collapse” consecutive sum nodes or product nodes.
To ensure the exact and efficient computation of various probabilistic queries, including marginalization and moment calculations, we must impose structural constraints on the circuit. Specifically, smoothness and decomposability (Peharz et al.,, 2015) are a set of sufficient conditions that ensure tractable computation of marginal and conditional probabilities. This tractability arises because smooth and decomposable circuits represent multilinear functions, which are known to support efficient marginalization (Broadrick et al.,, 2024). We provide details in Appendix A.
PCs can be viewed as latent variable models with discrete latent spaces (Peharz et al.,, 2016). Each sum node can be interpreted as introducing a discrete latent variable $Z$ that selects among its child distributions. Specifically, assigning $Z=i$ corresponds to choosing the $i$-th child of the sum node. By aggregating all such latent variables, the PC can be seen as defining a hierarchical latent variable model.
@@ 2.2 Expectation-Maximization
Expectation-Maximization (EM) is a well-known algorithm to maximize the log-likelihood given data $\boldsymbol{x}$ of a distribution defined by a latent variable model. Specifically, the distribution ${p}_{\boldsymbol{\phi}}(\mathbf{X})$ with parameters $\boldsymbol{\phi}$ is defined as $\sum_{\boldsymbol{z}}{p}_{\boldsymbol{\phi}}(\mathbf{X},\boldsymbol{z})$ over latents $\mathbf{Z}$. Our goal is to maximize
@@
EM can effectively maximize the above objective when ${p}_{\boldsymbol{\phi}}(\mathbf{X},\mathbf{Z})$ permits much simpler (or even closed-form) maximum likelihood estimation. It optimizes $\mathtt{LL}(\boldsymbol{\phi})$ by maximizing the following surrogate objective:
@@
EM updates the current parameters $\boldsymbol{\phi}$ by solving for $\boldsymbol{\phi}^{\prime}$ that maximizes $Q_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$, which is guaranteed to be a lower bound of $\mathtt{LL}(\boldsymbol{\phi}^{\prime})$ since
@@ 3 EM for Probabilistic Circuits
While variants of the EM algorithm have been proposed for training PCs in various contexts (Poon and Domingos,, 2011; Peharz,, 2015), their connection to the general EM objective $Q_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$ (cf. Eq. (3)) remains implicit. The lack of a unified formulation makes it difficult to fully understand the existing optimization procedures or to extend them to new settings, such as training with mini-batches of data, which is critical for scaling the optimizer to large datasets.
@@ 3.1 Revisiting Full-Batch EM
Given a PC ${p}_{\boldsymbol{\phi}}$ with log-parameters $\boldsymbol{\phi}$ (cf. Def. 1) and a dataset $\mathcal{D}$, $Q_{\boldsymbol{\phi}}^{\mathcal{D}}(\boldsymbol{\phi}^{\prime})$ equals the following up to a constant term irrelevant to $\boldsymbol{\phi}^{\prime}$:
$\frac{1}{|\mathcal{D}|}\sum_{\boldsymbol{x}\in\mathcal{D}}\left\langle\frac{\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})}{\partial\boldsymbol{\phi}},\boldsymbol{\phi}^{\prime}\right\rangle-\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime}),$ (4)
@@
The proof follows Kunstner et al., (2021) and is provided in Section B.1. Proposition 1 reveals that the EM update can be interpreted as maximizing a regularized first-order approximation of the log-likelihood. To see this, we rewrite Equation 4 by adding terms irrelevant to $\boldsymbol{\phi}^{\prime}$:
$\frac{1}{|\mathcal{D}|}\sum_{\boldsymbol{x}\in\mathcal{D}}\underbrace{\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})+\left\langle\frac{\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})}{\partial\boldsymbol{\phi}},\boldsymbol{\phi}^{\prime}-\boldsymbol{\phi}\right\rangle}_{\mathtt{LinLL}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(\boldsymbol{\phi}^{\prime})}-\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime}).$
In the above equation, the term $\mathtt{LinLL}_{\boldsymbol{\phi}}^{\boldsymbol{x}}(\boldsymbol{\phi}^{\prime})$ corresponds to the linearization of $\log{p}_{\boldsymbol{\phi}^{\prime}}(\boldsymbol{x})$ around the current parameters $\boldsymbol{\phi}$, capturing the local sensitivity of the log-likelihood to parameter changes. The KL term, $\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$, acts as a regularizer that penalizes large deviations in the joint distribution over $\mathbf{X}$ and $\mathbf{Z}$.
According to Proposition 1, solving for the updated parameters $\boldsymbol{\phi}^{\prime}$ requires computing two key quantities: the gradient of the log-likelihood $\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})/\partial\boldsymbol{\phi}$, and the KL divergence $\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$. To express these terms in closed form, we introduce the concept of top-down probabilities, which is defined by Dang et al., (2022).
@@
Given a PC p{p} parameterized by $\boldsymbol{\phi}$, we define the top-down probability $\mathtt{TD}(n)$ of a node $n$ recursively from the root node to input nodes:
[…]
@@
Assume the distributions defined by all nodes in a PC are normalized. For every $\boldsymbol{x}$, we have:
(i) $\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})/\partial\boldsymbol{\phi}=\partial\log\tilde{{p}}_{\boldsymbol{\phi}}(\boldsymbol{x})/\partial\boldsymbol{\phi}-\mathtt{TD}(\boldsymbol{\phi}),$
@@
The assumption that the PC is normalized is mild and practical. In Section 4 and Appendix C, we introduce a simple and efficient algorithm that adjusts the PC parameters to ensure normalization without affecting the structure of the circuit. We can now substitute the closed-form expressions for the gradient and the KL divergence into the general EM objective $Q_{\boldsymbol{\phi}}^{\mathcal{D}}(\boldsymbol{\phi}^{\prime})$, which converts the problem into [^5] We drop all terms that are independent of $\boldsymbol{\phi}^{\prime}$.
$\left\langle\frac{1}{|\mathcal{D}|}\sum_{\boldsymbol{x}\in\mathcal{D}}\frac{\partial\log\tilde{{p}}_{\boldsymbol{\phi}}(\boldsymbol{x})}{\partial\boldsymbol{\phi}}-\cancel{\mathtt{TD}(\boldsymbol{\phi})},\boldsymbol{\phi}^{\prime}\right\rangle+\cancel{\langle\mathtt{TD}(\boldsymbol{\phi}),\boldsymbol{\phi}^{\prime}\rangle}.$
@@
[…]
While this full-batch EM algorithm in Equation 5 has been derived in prior work (Choi et al.,, 2021; Peharz,, 2015), we recover it here through Proposition 1. This paves the way for a principled mini-batch EM algorithm by generalizing the objective $Q_{\boldsymbol{\phi}}^{\mathcal{D}}(\boldsymbol{\phi}^{\prime})$, as shown in the next section.